# Hierarchical Long Video Understanding with Audiovisual Entity Cohesion and Agentic Search

### v1

@@ 4.3.2 Analysis on Six Categories
To further assess the effectiveness of our method across different query types, we compare Ours with Ours_clip in terms of both accuracy and the average number of reasoning iterations required to answer queries across all six categories. As shown in Figure 3, our approach consistently achieves higher accuracy with fewer iterations compared to the non-hierarchical baseline, with particularly notable gains in entity recognition and reasoning. This improvement is expected, as hierarchical indexing provides readily accessible information at multiple granularities, enabling the agent to retrieve relevant context more efficiently and thereby reducing reasoning steps. It is worth noting that the summarization category in LVBench often involves fine-grained, step-level queries (e.g., “What does Mandy do after she stands in front of the judges’ table?”), which inherently require multi-turn reasoning rather than simple scene summarization.
@@ 5 Conclusion
In this work, we tackled the challenges of long video understanding by proposing a unified framework that integrates offline hierarchical video indexing with agentic multi-granularity retrieval. Our approach organizes video content across four levels, global, scene, segment, and entity, while incorporating audiovisual entity cohesion to maintain semantic consistency over extended temporal spans. Extensive experiments on LVBench demonstrate that our method significantly outperforms state-of-the-art baselines in overall accuracy, highlighting the effectiveness of hierarchical indexing and multimodal entity integration. These findings underscore the potential of structured, multi-level representations for advancing long video comprehension, a promising direction for future research.
Appendix A Visualization of Entity Coherence by Auditory Clues
In this section, we further illustrate the role of speaker identity in entity tracking. Figure 5 presents two examples. In these cases, characters undergo dramatic appearance changes, such as clothing, viewpoint shifts, and occlusion, throughout the video, complicating the merging of fragmented descriptions into a single entity. Leveraging speaker coherence, however, enables successful consolidation of these descriptions.
Figure 5: Cases: speaker identity for entity tracking.
Appendix B Additional Case Study
@@
Figure 6: Cases with multiple reasoning steps.
Appendix C Prompt for the Planner

### v2

@@ 4.3.2 Analysis on Six Categories
Figure 5: Comparison of accuracy and efficiency on six categories. “Acc” and “iter” denote accuracy and the average number of reasoning iterations, respectively.
To further assess the effectiveness of our method across different query types, we compare Ours with Ours_clip in terms of accuracy and the average number of reasoning iterations required to answer queries across all six categories. As shown in Figure 5, our approach consistently achieves higher accuracy with fewer iterations compared to the non-hierarchical baseline, with particularly notable gains in entity recognition and reasoning. This improvement is expected, as hierarchical indexing provides readily accessible information at multiple granularities, enabling the agent to retrieve relevant context more efficiently and thereby reducing reasoning steps. It is worth noting that the summarization category in LVBench often involves fine-grained, step-level queries (e.g., “What does Mandy do after she stands in front of the judges’ table?”), which require multi-turn reasoning rather than simple scene summarization.
@@ 5 Conclusion
In this work, we address the challenges of long video understanding with a unified framework that combines offline hierarchical video indexing and agentic multi-granularity retrieval. Our approach organizes video content into four levels, global, scene, segment, and entity, while incorporating audiovisual entity cohesion to maintain semantic consistency over extended temporal spans. Extensive experiments show that our method consistently outperforms state-of-the-art baselines, demonstrating the effectiveness of hierarchical indexing and audiovisual entity consolidation. These results highlight the promise of structured, multi-level representations for advancing long video comprehension and motivate further research in this direction.
Supplementary Material
The supplemental material contains additional implementation details as well as more results and discussions.
Appendix A Entity-Centric Re-Captioning
After consolidation, each canonical entity $\tilde{e}_{j}$ is associated with a global description and a set of linked segments $Q_{j}$. During retrieval, incorporating all linked segments of top-$K_{1}$ entities can be computationally expensive and may introduce query-irrelevant noise. Such noise can weaken embedding-based matching; while LLM-based re-ranking may mitigate this issue, it incurs higher computational cost. To address this, we further introduce an entity-centric re-captioning process during offline construction. For each linked segment $i$ and entity $\tilde{e}_{j}$, we generate a focused description $\tilde{C}^{t}_{i,j}$ using an LLM that summarizes the entity’s appearance, actions, and events within the segment, while excluding irrelevant context. The final entity database contains both canonical entities and fine-grained entity-segment descriptions: $\tilde{\mathcal{E}}=\{\tilde{\mathcal{E}}_{g}\,;\,\tilde{\mathcal{E}}_{e}\}$, where $\tilde{\mathcal{E}}_{e}=\{\tilde{C}^{t}_{i,j}|j=1,2,\ldots,J,i\in Q_{j}\}$. During retrieval, we first match entities in the embedding space of $\tilde{\mathcal{E}}_{g}$, then re-rank linked segments using similarity between the query and $\tilde{C}^{t}_{i,j}$, selecting the top-$K_{2}$ segments for precise grounding. This design balances retrieval precision and computational efficiency, avoiding excessive LLM overhead. In our implementation, $K_{1}$ and $K_{2}$ are set to 20 and 16, respectively.
Appendix B Multi-Granularity Tools
Here we present details of the whole tool set denoted by
Global Scene Browse. This tool $T_{\text{scene}}$ supports coarse-grained navigation and scene localization along the video timeline. Given a user query $q$ and the scene collection $D=\tilde{\mathcal{S}}$, it identifies and summarizes the most relevant scenes with an LLM, returning their storyline and corresponding timestamps $\tau^{\prime}$. The agent tends to invoke this tool for complex or ambiguous queries involving multiple events or temporal dependencies.
Segment Caption Search. This tool $T_{\text{caption}}$ performs fine-grained text-based retrieval within specified temporal ranges. Given the user query $q$ and the segment database $D=\tilde{\mathcal{C}}$, the tool retrieves the most semantically relevant segment descriptions $r$ along with their associated time spans $\tau^{\prime}$. This is achieved through cosine similarity matching between the query embedding and pre-computed caption embeddings for all video segments, ensuring efficient and accurate retrieval of localized content.
Segment Visual Search. To capture visual cues that may be overlooked in textual descriptions, the Segment Visual Search tool $T_{\text{visual}}$ complements $T_{\text{caption}}$. While the latter relies on text embeddings, $T_{\text{visual}}$ leverages cross-modal embeddings generated by the UNITE framework [11]. This design enables retrieval driven by rich visual semantics aligned with the query, ensuring that visually salient details are incorporated into the search process.
Entity Search. This tool $T_{\text{entity}}$ supports high-level, entity-centric retrieval across large temporal ranges. Given an entity-related query $q$, the tool first retrieves the top-$K_{1}$ most relevant entities from database $\tilde{\mathcal{E}_{g}}$ based on their descriptions in the pre-computed embedding space. For all segments linked to these entities, it then performs a second-stage reranking to select the top-$K_{2}$ most relevant segments from the entity-centric database $\tilde{\mathcal{E}_{e}}$, using the same query. Finally, the tool applies entity-aware re-captioning to the retrieved segments, generating a coherent response $r$ enriched with precise timestamps $\tau^{\prime}$.
Inspection Tool. This tool $T_{\text{inspect}}$ provides fine-grained temporal inspection to support detailed reasoning. It consists of two complementary modules: Clip Caption Inspect ($T_{\text{inspect}}^{\text{tex}}$) and Visual Inspect ($T_{\text{inspect}}^{\text{vis}}$). The $T_{\text{inspect}}^{\text{tex}}$ examines coarse textual descriptions to determine what occurred during a specified time span. For example, for a query such as “What does the protagonist do after he jumps down the stairs?”, the tool inspects subsequent time ranges to identify the protagonist’s next actions after locating the event “he jumps down the stairs” in previous iteration. The $T_{\text{inspect}}^{\text{vis}}$ tool leverages a VLM to perform precise visual verification within a given time range. Due to frame limits of the VLM, this inspection focuses on short intervals, ensuring accurate visual grounding for fine-grained queries.
Figure 6: Cases with multiple reasoning steps.
Appendix C Other Technical Details
We leverage agglomerative clustering with a threshold of 0.4 for entity clustering stage. A chunk size of 24 with an overlap of 3 is employed for scene segmentation. Due to the API frame limit, the Ours (2 fps) variant generates captions for two 15-second sub-segments, each sampled with 30 frames, and concatenates them into a single description.
Appendix D Additional Case Study
@@
Appendix E Efficiency
Our semantic-consistent hierarchy enables more efficient navigation, achieving higher accuracy with fewer reasoning steps and less runtime compared with DVD, as shown in Table 4.
Table 4: Comparison of average number of iterations and runtime (second per query).
Methods Iteration Runtime (s)
DVD 7.6 151.0
Ours 4.2 98.7
Appendix F Proprietary models
Our API versions are GPT-4.1 (2025-04-14) and o3 (2025-04-16). Variance of three runs on LVBench is 0.149, which demonstrates the robustness of our method. Furthermore, Our method achieves an competitive accuracy of 75.8% on LVBench using open-source models (DeepSeek-R1-0528 for reasoning + Qwen3-VL-32B-Instruct for visual inspection).
Appendix G Prompt for the Planner